OpenStudy (anonymous):

the length of a rectangle is 7 meters longer than its width. what is the width of this rectangle if its perimeter is equal to 86 meters?

OpenStudy (whpalmer4):

\[A = lw\]\[P=2l+2w\]\[l = 7+w\] Substitute the 3rd equation into the second, set the result equal to the known perimeter, and solve for the length. Then use the third equation to find the width knowing the length, and plug both width and length into the first equation to find the area.

OpenStudy (texaschic101):

P = 2L + 2W P = 86 L = W + 7 now just sub all that into your formula and solve for W (width)